# Wonderful Team: Zero-Shot Physical Task Planning with Visual LLMs

### v5

@@ D.1 Understanding What Each Part of Wonderful Team Does
Appendix E Ablation Studies: VLLMs’ Spatial Reasoning Limitations and Potentials
E.1 Evaluating VLLM’s Spatial Understanding
@@ E.1.3 Deviation Analysis
E.2 Evaluating VLLMs’ Error Recognition and Correction
@@
E.2.1 Experimental Setup
To test this ability, we randomly generated 4 types of bounding boxes around the circle of interest. Examples are shown in 23. The types are:
@@ E.2.2 Results
Appendix F Comparison with Other Methods
F.1 Replicating Natural Language as Policies Using gpt-4o
@@
F.2 Comparison with PIVOT
@@
  • PIVOT identifies a single action point for the target object, maintaining this as the goal throughout their iterative process. In contrast, our method offers two distinct workflows that the grounding manager can choose from before localization. When selecting an area point, such as a position between a box and a frame, we also employ point selection. However, for object selection, our method first identifies a center point, then determines a bounding box of appropriate size, and iteratively refines this bounding box until it is accurate. The grounding manager then selects an actionable point within the bounded area. We found that this bounding box process greatly enhances robustness and precision, especially for smaller objects or manipulation tasks that require more precise control. We further ablate and discuss this in Appendix F.2.
@@
In Table 17, we compare the experimental results of our method with those from PIVOT. While PIVOT originally utilizes GPT-4V in its framework, we implemented their approach using the more advanced GPT-4O to ensure a fair comparison. Our replication of their framework was carried out to the best of our knowledge to highlight the differences and performance improvements. Additionally, we include results obtained from their official HuggingFace demo to demonstrate the performance of their original implementation. For example output of different grounding approaches, please see 31.
@@
Parallel Runs: The original study also employs a parallel call strategy. To combine results from different runs, they explored two approaches: (1) fitting a new action distribution from the output actions and returning it, and (2) selecting a single best action using a VLM query. In our implementation, we used the second approach with “3 Iterations 3 Parallel“ combinations to enhance robustness in our comparison. Additionally, while the original implementation uses the same sampling radius for both width and height, we addressed this by defining separate radii for the shorter and longer edges of the input image. Grounding Team Only: Since PIVOT’s framework is primarily comparable to our grounding team, which focuses on processing object descriptions rather than broader tasks, we isolated the grounding component for a direct comparison with their method. Success Evaluation: For evaluation, we conducted 10 runs on different objects from a set of varied initial frames. A task was considered successful if the center point label of each target object had at least half of its area within the object’s boundary or if the center point fell within a specific range around the target area center, ensuring successful picking. Refer to caption Figure 30: Screenshot of HuggingFace PIVOT Demo Failure Mode Discussions It’s notable that PIVOT’s output on tabletop tasks does not over-perform the direct output from GPT-4o. However, this is with the help of the labeled coordinate system, which significantly enhances precision in quantification, as discussed in our motivation section. We further discuss the possible explanations of PIVOT failures: Incomplete Sampling Coverage: In 30, when attempting to select the left object, the initial sampling failed to provide sufficient coverage, with the majority of points being sampled from the center of the image and scattering on the purple paisley letter “V” instead of the target object with blue and purple stripes. As a result, subsequent iterations were confined to a suboptimal region, ultimately leading to poor final results. Difficulty in Recovery: During our implementation, we identified a critical limitation in the sampling strategy: if the sampling radius is too small, it becomes difficult to recover from an inadequate initial selection. Conversely, if the sampling radius is too large, the framework struggles to converge, as the sampled actions may scatter too broadly, reducing the effectiveness of the refinement process. Lack of Iterative Continuity: Another factor that may explain PIVOT’s low performance in precise location finding is the lack of continuity between iterations. Although the new set of actions is sampled from a distribution fitted using previously selected promising actions, there is a notable discontinuity in the process. For instance, if a good point is identified during one iteration, it is not guaranteed to be preserved in subsequent iterations. The framework’s fixed number of resampling processes means it cannot exit the process once a good point is found, potentially resulting in the loss of successful actions. This resampling process can lead to promising actions being either diluted or completely discarded in the next round due to inherent randomness, causing inefficiencies and inconsistencies as the framework may fail to build on previous successes. Messy Annotations: Additionally, the framework’s annotations can become cluttered, leading to a loss of crucial information from the original image. Unlike our approach, which maintains a clear connection to the original image to preserve full context, PIVOT’s method can lose track of the overall scene, making it difficult to refine action points effectively. This loss of context can be particularly detrimental in scenarios where precision and consistency are critical. Refer to caption Figure 31: Example Outputs - Wonderful Team vs PIVOT Point Selection vs. Bounding Box: Since the PIVOT method is inherently more similar to our area/point approach discussed earlier—where points are selected throughout the process without the aid of bounding boxes—we further compare PIVOT’s outputs with both our bounding box approach and our point approach. Figure 31 provides insight into how these methods perform relative to each other. While both PIVOT and our area/point approach can get reasonably close to the desired objects, they often lack the precision required for tasks involving small objects or when execution demands more accuracy than just proximity to the object. In Figure 32, we present example executions using the results from these methods. The task involves stacking the purple and blue striped letter “V” on top of the blue letter “V,“ followed by stacking the purple paisley letter “V” on top. For this execution, we used the PIVOT results from our implementation using gpt-4o, as the HuggingFace outputs were less reliable, with all points concentrated on the same object. The execution screenshots reveal that points not accurately placed on the object lead to failures in picking it up. On the bottom row of Figure 32, even though both points for the first pick-and-place action are technically correct, the misalignment causes the stacking task to partially fail, as the letters “V“ are not properly aligned, resulting in an unsuccessful stack. These results highlight the importance of considering whether a bounding box is needed in the iterative process. With the current level of visual reasoning skills in models, we found that incorporating a bounding box significantly enhances precision, reduces hallucinations, and adds robustness to the execution. Refer to caption Figure 32: Example Executions - Wonderful Team vs PIVOT
@@
F.3 Comparison with Language Models as Zero-Shot Trajectory Generators
F.3.1 Key Differences
@@
F.3.2 Single Agent vs Multi-Agent
@@
F.3.3 Benefits of Using a Multi-Agent System
@@
F.3.4 Experimental Comparison in Fetch
@@
F.4 Comparison with MOKA
F.4.1 Key Differences
@@ 1. Vision-Language Integration:
  • Like Trajectory Generator, MOKA relies on separate vision models (GroundingDINO + SAM, comparable to LangSAM). As discussed in Appendix F.3, these models struggle with detecting objects that are context-specific, complex, or ambiguously described.
@@ 2. System Architecture:
  • MOKA: Uses annotated images from GroundingDINO + SAM, similar to PIVOT (Appendix F.2). The VLLM selects a point from these annotations, but struggles with scattered or unreadable points for small or ambiguous objects, as illustrated in Figure 36.
@@ F.4.2 Error Analysis
Similar to Trajectory Generator’s pipeline—which fails on complex objects and longer-horizon planning and reasoning (details discussed in Appendix F.3)—MOKA also suffers from object detection errors due to the separation of planning and grounding. However, while Trajectory Generator consistently generates complete plans without execution errors, half of MOKA’s failure cases result in incomplete plans. These failures are caused by code crashes, which prevent the generation of any plan.
@@
  • Vision Model Limitations: GroundingDINO, like other vision models such as OWL-ViT and LangSAM, struggles in complex environments or with less common objects. These limitations, extensively discussed in Appendix F.3, result in incomplete or failed detections that hinder the overall pipeline. For instance, intricate object arrangements, subtle textures, or unusual configurations exacerbate detection challenges, further limiting the model’s reliability.

### v6

@@ E.1 Understanding What Each Part of Wonderful Team Does
Appendix F Ablation Studies: VLLMs’ Spatial Reasoning Limitations and Potentials
F.1 Evaluating VLLM’s Spatial Understanding
@@ F.1.3 Deviation Analysis
F.2 Evaluating VLLMs’ Error Recognition and Correction
@@
F.2.1 Experimental Setup
To test this ability, we randomly generated 4 types of bounding boxes around the circle of interest. Examples are shown in 29. The types are:
@@ F.2.2 Results
Appendix G Comparison with Other Methods
G.1 Replicating Natural Language as Policies Using GPT-4o
@@
G.2 Comparison with PIVOT
@@
  • PIVOT identifies a single action point for the target object, maintaining this as the goal throughout their iterative process. In contrast, our method offers two distinct workflows that the grounding manager can choose from before localization. When selecting an area point, such as a position between a box and a frame, we also employ point selection. However, for object selection, our method first identifies a center point, then determines a bounding box of appropriate size, and iteratively refines this bounding box until it is accurate. The grounding manager then selects an actionable point within the bounded area. We found that this bounding box process greatly enhances robustness and precision, especially for smaller objects or manipulation tasks that require more precise control. We further ablate and discuss this in Appendix G.2.
@@
In Table 17, we compare the experimental results of our method with those from PIVOT. While PIVOT originally utilizes GPT-4V in its framework, we implemented their approach using the more advanced GPT-4o to ensure a fair comparison. Our replication of their framework was carried out to the best of our knowledge to highlight the differences and performance improvements. Additionally, we include results obtained from their official HuggingFace demo to demonstrate the performance of their original implementation. For example output of different grounding approaches, please see 37.
@@
Parallel Runs: The original study also employs a parallel call strategy. To combine results from different runs, they explored two approaches: (1) fitting a new action distribution from the output actions and returning it, and (2) selecting a single best action using a VLM query. In our implementation, we used the second approach with “3 Iterations 3 Parallel“ combinations to enhance robustness in our comparison. Additionally, while the original implementation uses the same sampling radius for both width and height, we addressed this by defining separate radii for the shorter and longer edges of the input image. Grounding Team Only: Since PIVOT’s framework is primarily comparable to our grounding team, which focuses on processing object descriptions rather than broader tasks, we isolated the grounding component for a direct comparison with their method. Success Evaluation: For evaluation, we conducted 10 runs on different objects from a set of varied initial frames. A task was considered successful if the center point label of each target object had at least half of its area within the object’s boundary or if the center point fell within a specific range around the target area center, ensuring successful picking. Refer to caption Figure 36: Screenshot of HuggingFace PIVOT Demo Failure Mode Discussions It’s notable that PIVOT’s output on tabletop tasks does not over-perform the direct output from GPT-4o. However, this is with the help of the labeled coordinate system, which significantly enhances precision in quantification, as discussed in our motivation section. We further discuss the possible explanations of PIVOT failures: Incomplete Sampling Coverage: In 36, when attempting to select the left object, the initial sampling failed to provide sufficient coverage, with the majority of points being sampled from the center of the image and scattering on the purple paisley letter “V” instead of the target object with blue and purple stripes. As a result, subsequent iterations were confined to a suboptimal region, ultimately leading to poor final results. Difficulty in Recovery: During our implementation, we identified a critical limitation in the sampling strategy: if the sampling radius is too small, it becomes difficult to recover from an inadequate initial selection. Conversely, if the sampling radius is too large, the framework struggles to converge, as the sampled actions may scatter too broadly, reducing the effectiveness of the refinement process. Lack of Iterative Continuity: Another factor that may explain PIVOT’s low performance in precise location finding is the lack of continuity between iterations. Although the new set of actions is sampled from a distribution fitted using previously selected promising actions, there is a notable discontinuity in the process. For instance, if a good point is identified during one iteration, it is not guaranteed to be preserved in subsequent iterations. The framework’s fixed number of resampling processes means it cannot exit the process once a good point is found, potentially resulting in the loss of successful actions. This resampling process can lead to promising actions being either diluted or completely discarded in the next round due to inherent randomness, causing inefficiencies and inconsistencies as the framework may fail to build on previous successes. Messy Annotations: Additionally, the framework’s annotations can become cluttered, leading to a loss of crucial information from the original image. Unlike our approach, which maintains a clear connection to the original image to preserve full context, PIVOT’s method can lose track of the overall scene, making it difficult to refine action points effectively. This loss of context can be particularly detrimental in scenarios where precision and consistency are critical. Refer to caption Figure 37: Example Outputs - Wonderful Team vs PIVOT Point Selection vs. Bounding Box: Since the PIVOT method is inherently more similar to our area/point approach discussed earlier—where points are selected throughout the process without the aid of bounding boxes—we further compare PIVOT’s outputs with both our bounding box approach and our point approach. Figure 37 provides insight into how these methods perform relative to each other. While both PIVOT and our area/point approach can get reasonably close to the desired objects, they often lack the precision required for tasks involving small objects or when execution demands more accuracy than just proximity to the object. In Figure 38, we present example executions using the results from these methods. The task involves stacking the purple and blue striped letter “V” on top of the blue letter “V,“ followed by stacking the purple paisley letter “V” on top. For this execution, we used the PIVOT results from our implementation using GPT-4o, as the HuggingFace outputs were less reliable, with all points concentrated on the same object. The execution screenshots reveal that points not accurately placed on the object lead to failures in picking it up. On the bottom row of Figure 38, even though both points for the first pick-and-place action are technically correct, the misalignment causes the stacking task to partially fail, as the letters “V“ are not properly aligned, resulting in an unsuccessful stack. These results highlight the importance of considering whether a bounding box is needed in the iterative process. With the current level of visual reasoning skills in models, we found that incorporating a bounding box significantly enhances precision, reduces hallucinations, and adds robustness to the execution. Refer to caption Figure 38: Example Executions - Wonderful Team vs PIVOT
@@
G.3 Comparison with Language Models as Zero-Shot Trajectory Generators
G.3.1 Key Differences
@@
G.3.2 Single Agent vs Multi-Agent
@@
G.3.3 Benefits of Using a Multi-Agent System
@@
G.3.4 Experimental Comparison in Fetch
@@
G.4 Comparison with MOKA
G.4.1 Key Differences
@@ 1. Vision-Language Integration:
  • Like Trajectory Generator, MOKA relies on separate vision models (GroundingDINO + SAM, comparable to LangSAM). As discussed in Appendix G.3, these models struggle with detecting objects that are context-specific, complex, or ambiguously described.
@@ 2. System Architecture:
  • MOKA: Uses annotated images from GroundingDINO + SAM, similar to PIVOT (Appendix G.2). The VLLM selects a point from these annotations, but struggles with scattered or unreadable points for small or ambiguous objects, as illustrated in Figure 42.
@@ G.4.2 Error Analysis
Similar to Trajectory Generator’s pipeline—which fails on complex objects and longer-horizon planning and reasoning (details discussed in Appendix G.3)—MOKA also suffers from object detection errors due to the separation of planning and grounding. However, while Trajectory Generator consistently generates complete plans without execution errors, half of MOKA’s failure cases result in incomplete plans. These failures are caused by code crashes, which prevent the generation of any plan.
@@
  • Vision Model Limitations: GroundingDINO, like other vision models such as OWL-ViT and LangSAM, struggles in complex environments or with less common objects. These limitations, extensively discussed in Appendix G.3, result in incomplete or failed detections that hinder the overall pipeline. For instance, intricate object arrangements, subtle textures, or unusual configurations exacerbate detection challenges, further limiting the model’s reliability.
G.5 Comparison with Fine-Tuned VLA Models
Our main experiments (Section 5) focus on methods that, like ours, employ pretrained large language models with vision capabilities (e.g., GPT-4o) without additional fine-tuning on robotics data. In contrast, another branch of work trains Vision-Language-Action (VLA) models on large-scale robot-manipulation datasets (see Section 4). Here, we evaluate two open-sourced state-of-the-art VLA systems:
  • RT-1 (Brohan et al., 2022): A variant of Google DeepMind’s RT-1 family. RT-1 is a Transformer-based model that processes sequences of images and natural-language instructions to produce real-time robotic actions, fine-tuned on large-scale robotics data.
  • OpenVLA (Kim et al., 2024): A 7B-parameter vision-language-action model combining LLaMA 2 with a fused visual encoder (DINOv2 + SigLIP features) and an action-detokenizer. Fine-tuned variants exist for WidowX and Google robots from Open X-Embodiment, plus a version trained on the simulated Libero environment (Liu et al., 2023a).
G.5.1 Experimental Setup
Environment and Robot Choice.
Directly evaluating fine-tuned VLA models like RT-1 and OpenVLA on the same tasks as our framework is challenging due to their fine-tuned nature and inherent dependencies:
  1. 1.
    Robot-Specific Action Output: RT-1 and OpenVLA are designed to generate actions tailored to specific robots (e.g., Google robots or WidowX). Transferring these models to a robot available in our lab would require additional fine-tuning, which itself demands significant data collection and training efforts. This makes a direct comparison infeasible without introducing additional biases.
  2. 2.
    Sensitivity to Distribution Shifts. Fine-tuned models like RT-1 and OpenVLA rely on datasets with specific assumptions about backgrounds, object layouts, and robot kinematics. Even small variations, such as changes in camera pose, can cause significant performance drops—up to 50% in both simulation and real-world settings for RT-1, as reported by Li et al. (2024).
(a) Real
(b) SIMPLER simulation
(c) Performance Correlation
Figure 48: Figures adapted from Li et al. (2024). (a) Real-world setup mimicking Open X Embodiment videos, (b) the corresponding high-fidelity SIMPLER simulation, and (c) strong correlation between SIMPLER and real-world task performance.
To make the comparison as fair as possible, we leverage SIMPLER (Li et al., 2024), a real-to-sim framework shown by Li et al. (2024) to correlate strongly with real-world performance (Figure 48(c)). By reconstructing our tasks in SIMPLER Environment, we hope to reduce the potential domain gap for RT-1 and OpenVLA evaluation. Nevertheless, discussion threads (e.g., GitHub issues) indicate that OpenVLA may be more sensitive to real-to-sim transfer than RT-based models, so we also test an OpenVLA variant trained on the Libero environment (Liu et al., 2023a).
Task Choice.
We do not test all tasks from our suite because:
  • Lack of multimodal prompt support: Many tasks in VIMABench rely on multimodal inputs, such as image-based prompts that reference objects or actions directly. However, RT-1 and OpenVLA lack the ability to process integrated text-image prompts (e.g., as supported by GPT-4o), accepting only text instructions and environmental frames. This limitation makes certain tasks incompatible with these models.
  • Challenges reconstructing complex tasks in simulated environment: Fine-grained tasks such as opening a bottle cap or drawing a star require precise simulation physics, which are not fully supported by current physics engines. These tasks would likely introduce additional noise unrelated to the core evaluation of the models.
  • Early failures on simpler tasks. Preliminary evaluations showed that both RT-1 and OpenVLA consistently failed on basic pick-and-place scenarios, with near 0% success rates in many cases. Given these results, we prioritize simpler tasks to pinpoint where and why these models fail, without introducing unnecessary complexity.
Rather than performing an exhaustive comparison across all tasks, our goal is to explore three critical aspects of fine-tuned VLA model performance:
  1. 1.
    Can these models understand implicit prompts, such as those in our implicit-goal tasks?
  2. 2.
    Are they capable of handling long-horizon manipulations involving multiple sequential substeps?
  3. 3.
    How well do they generalize to uncommon objects or patterns, such as those encountered in VIMABench but potentially out-of-distribution for their training data?
To address these questions, we specifically test two tasks (and their variations):
  1. 1.
    Fruit Placement: From our implicit-goal suite, designed to test reasoning and implicit planning.
  2. 2.
    Simple Object Manipulation: A VIMABench-inspired pick-and-place scenario representing the simplest form of long-horizon tasks.
(a) Fruit Placement Setup in SIMPLER
(b) Simpler Object Manipulation Setup in SIMPLER
Figure 49: Evaluation setups for tasks in the SIMPLER environment. (a) Fruit Placement task setup: objects are placed in color-matched regions based on implicit goal inference. (b) Simple Object Manipulation task setup: basic pick-and-place tasks adapted from VIMABench. Both setups replicate real-world configurations with high fidelity, ensuring consistency in object placement, camera angles, and environmental details for fair comparisons across methods.
We provide a Colab notebook [^1] https://colab.research.google.com/drive/1C9JW4MXLNyN0LmqcrOqNEm7C6ghgfB9Z#scrollTo=H81NNh_jdIgP for reproducing these setups and evaluating both RT-1 and OpenVLA.
G.5.2 Experimental Results and Ablations
Overall Success on Two Tasks.
Despite efforts to minimize the domain gap (Figures 49(a) and 49(b)), RT-1 and OpenVLA achieve limited success on the evaluated tasks in SIMPLER Environment. Specifically, on the Fruit Placement task and the simplest VIMABench pick-and-place scenario, RT-1 yields only a 10% success rate on the latter, while OpenVLA fails to complete either task, achieving 0% success in both cases. In contrast, our Wonderful Team framework achieves a perfect 100% success rate across both tasks (Table 20).
Task Wonderful Team RT-1 OpenVLA
Fruit Placement (Implicit) 100% 0% 0%
Simple Object Manipulation (VIMABench) 100% 10% 0%
Table 20: Success rates (%) for Wonderful Team, RT-1, and OpenVLA on two simplified tasks in SIMPLEREnv. Each experiment was conducted over 10 trials with a maximum of 300 steps per trial.
Below, we further break down why these failures occur using ablation experiments.
G.5.3 Ablation Study: Fruit-Placement Task Variations
The original fruit-placement task, described in Appendix D.3, features a setup with three colored papers on a table, three to four fruits, and the potential for pre-misplaced fruits that require additional reasoning to identify and correctly re-place them. The task prompt is:
This setup tests implicit reasoning, such as inferring relationships between objects and spatial reasoning for proper alignment with the colored areas (Figure 21(a))
To analyze the performance of RT-1 and OpenVLA in handling progressively complex scenarios, we conducted an ablation study using the Fruit Placement Task. The setup is illustrated in Figure 50, where complexity is incrementally added by introducing more fruits, colors, and reasoning challenges.
Figure 50: Ablation study setups for the Fruit Placement Task. Starting from the simplest configuration, complexity is progressively added.
Table 21 summarizes the grasping and task completion success rates for RT-1 across various levels of task complexity. OpenVLA consistently achieved a 0% success rate across all configurations, including the simplest setup. A common failure mode observed for OpenVLA was the robot remaining stationary throughout all 300 steps, unable to initiate any actions. This aligns with prior community observations, which highlight OpenVLA’s significant sensitivity to distribution shifts.
RT-1 performs well in simpler configurations, successfully grasping the apple and placing it in the red area under minimal scene complexity. However, as the task becomes more challenging, such as introducing distractors or implicit reasoning prompts, RT-1’s performance declines significantly.
Colors Fruits Prompt RT-1 Grasped (%) RT-1 Success (%)
1 1 Place the apple in the red area. 100 70
1 2 Place the apple in the red area. Ignore the banana. 90 40
1 2 Place the apple in the red area. 70 20
1 2 Place each fruit in the area that matches its color, if such an area exists. 10 0
2 1 Place the apple in the red area. 100 30
2 1 Place each fruit in the area that matches its color, if such an area exists. 0 0
2 2 Place the apple in the red area. 70 30
2 2 Place each fruit in the area that matches its color, if such an area exists. 10 0
Table 21: Ablation results for Fruit Placement Task variations. RT-1 success rates (%) for grasping the correct object and completing the task are shown as complexity increases. OpenVLA results are omitted as all trials resulted in 0% success.
G.5.4 Ablation Study: VIMABench Simple Manipulation Task Variations
Figure 51: Ablation study on VIMABench focusing on pattern changes. Success rates for grasping vary significantly depending on the object pattern.
The VIMABench ablation study primarily examines the impact of pattern changes on task performance. As shown in Figure 51, RT-1’s success rate for grasping dropped from 90% for a red circle to 40% for a purple paisley circle. This shows the model’s sensitivity to less common object or pattern, a limitation that can hinder its robustness in scenarios with diverse or out-of-distribution objects and patterns.
Additional Evaluation of OpenVLA in Libero.
Given OpenVLA’s consistent 0% success rate in the SIMPLER Environment across tasks, we conducted additional evaluations using Libero (Liu et al., 2023a), a dataset specifically used to fine-tune OpenVLA-Libero. For in-distribution Libero objects, which primarily resemble sauce bottles and other common objects, OpenVLA achieved a 70% success rate in simple pick-and-place tasks. However, when the objects were replaced with pure, vibrant-colored cylinders (e.g., red, green, and blue), as shown in Figure 52, the success rate dropped to 0%. This result suggests that the failures observed in the SIMPLER Environment were not solely due to real-to-sim transfer issues but were instead indicative of OpenVLA’s limited ability to generalize to out-of-distribution objects or layouts.
Figure 52: Evaluation of OpenVLA on Libero with a combination of in-distribution objects and out-of-distribution color-coded setups.
G.5.5 Discussion
Our results highlight the trade-offs between specialized fine-tuned VLA models and general-purpose VLLM-based methods.
Performance on novel and complex tasks. While RT-1 and OpenVLA excel at tasks aligned with their training distributions, they struggle significantly under even mild domain shifts, such as novel objects, unexpected layouts, or implicit multi-step reasoning. For example, OpenVLA often remained idle when confronted with unseen conditions, and RT-1’s performance deteriorated sharply as task complexity increased.
Advantages of fine-tuned VLA models. Fine-tuned VLA models remain valuable in scenarios closely aligned with their training data, where they can generate valid actions without additional post-processing. However, extending their capabilities to handle variations in object properties, camera viewpoints, and environmental contexts is a critical challenge.
The trade-off between reasoning and action generation. A shift toward action generation through fine-tuning raises another question: does this focus come at the expense of higher-level reasoning and planning abilities? Investigating this trade-off is essential for understanding the limitations and opportunities of fine-tuned models.
Generalist robotics systems and the roles of VLLMs and VLAs. This brings us to a key question for future research: should a single fine-tuned VLA model manage the entire robotics pipeline—from perception to planning to action—or is a modular, hierarchical approach more effective? The trade-off between reasoning and action generation lies at the core of this discussion, as a greater focus on action outputs may come at the expense of reasoning depth and planning flexibility.
From our experiments with Wonderful Team, even with GPT-4o—a state-of-the-art general-purpose reasoning model—a hierarchical structure with self-correction mechanisms proved essential for achieving robust performance across diverse robotics tasks. This suggests that robotics systems benefit significantly from a division of responsibilities: VLA models can excel as specialized modules for precise, subgoal-level action generation, while VLLMs, with their superior reasoning and world-modeling capabilities, handle high-level planning, abstract decision-making, and failure correction through language. Importantly, our insights suggest that VLLMs and VLAs need not be substitutes for one another but rather complementary components.